OpenStudy (anonymous):

How to solve 4k^-2/3=9

OpenStudy (rulnick):

k = 8/27

OpenStudy (rulnick):

... because k^(-2/3) = 9/4 or k^(2/3) = 4/9.

OpenStudy (rulnick):

... need more? k = (4/9)^(3/2) or sqrt((4/9)^3).

OpenStudy (anonymous):

Ahh, I see what I did wrong

OpenStudy (anonymous):

No, I see it. thanks and this site is awesome!

OpenStudy (rulnick):

welcome.

hero (hero):

@Diyadiya :D

OpenStudy (anonymous):

Sorry about not giving you the credit. I appreciated your help!!

OpenStudy (rulnick):

no problem. it happens. thanks for checking back!